Try yourself 1 ► The owner of a small tailoring shop keeps careful records of all alterations. Two common alterations to menPlease help me solve this and how do I decide which confidence level I should use z or t?
Answer #1

Sol:

a)

we want to find the 95% confidence interval for the mean alteration of the inseams on men's pants.

X1 + Z2/2*01/Vn1

0.74 + 20.05/2 * 0.22/33

04-1.96 96 ¥00383 ; Z- value from std. normal table at alpha=0.05

04-0071

10.66.0.82

The sample mean are lies between 0.66 and 0.82.

b)

Therefore,

E=0.05

The sample size n is

n= Za/201 E

Zo_05/2 * 0.22、 0.05

  - 1.95 : 0.22 005 ; Z- value from std. normal table at alpha=0.05

  =74

The 74 large a sample is necessary foe 95 % confidence interval .

c)

we want to find the 90% confidence interval for the mean alteration of the waists of sports coasts.

X2 + Z2/2 * 02/Vn2

1.05\pm Z_{0.10/2}*0.37/\sqrt{42}

1.05\pm 1.64*0.0571 ; Z- value from std. normal table at alpha=0.10

1.05\pm0.0936

0.96, 1.14

The sample mean are lies between 0.96 and 1.14.

d).

Therefore,

E=0.07

The sample size n is

n= Za/202 E

20.10/2 * 0.37 0.07

  =(\frac{1.64*0.37}{0.07})^2 ; Z- value from std. normal table at alpha=0.10

  75

The 75 large a sample is necessary foe 90 % confidence interval .

If you Satisfy with Answer, Please give me "Thumb Up". It is very useful for me.
Thank you for your support.